OpenStudy (anonymous):

Find the slope of the line that passes through the points (2, 1) and (-4, -5). -1 1 correct? 7/5

OpenStudy (mertsj):

Subtract the y values. What do you get?

OpenStudy (mertsj):

Subtract the x values? What do you get?

OpenStudy (mertsj):

Put the first answer over the second number.

OpenStudy (anonymous):

huh? so its not 1? I know its not -1

OpenStudy (mertsj):

1-(-5)=6 2-(-4)=6 6/6=1